OpenStudy (anonymous):

HELP PLEASE!!! WILL MEDAL!!! Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4).

OpenStudy (anonymous):

@chosenmatt

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@demonchild99

OpenStudy (texaschic101):

first we need to find the slope of the line 4x - 3y = -12. The reason we do this is because a parallel line will have the same slope. Do you know how to find the slope of the equation ?

OpenStudy (anonymous):

Kind of

OpenStudy (texaschic101):

in y = mx + b form, the slope is in the m position. So lets put the equation in y = mx + b form and find the slope. 4x - 3y = -12 --- subtract 4x from both sides -3y = -4x - 12 --- now divide both sides by -3 y = 4/3x + 4 the number in the m position is 4/3...that is your slope now we use y = mx + b slope(m) = 4/3 (-3,4)...x = -3 and y = 4 now we sub and solve for b, the y intercept. 4 = 4/3(-3) + b 4 = -4 + b 4 + 4 = b 8 = b so your parallel equation is : y = 4/3x + 8

OpenStudy (texaschic101):

questions ?

OpenStudy (texaschic101):

what do you not understand ?

OpenStudy (anonymous):

Not really... thanks! Ill write this down so I can study this for tomorrow.. :)

OpenStudy (trojanpoem):

You understood ?